OpenStudy (anonymous):

6h * 4h --------- 2h can you explain how to get answer aswell..... :)

Mehek (mehek14):

you multiply 6 * 4 and divide by 2 keep the h

OpenStudy (kendricklamar2014):

\[6h * 4h = 24h^2\]

OpenStudy (kendricklamar2014):

\[\left(\begin{matrix}24h \\ 2\end{matrix}\right) = 12h\]

OpenStudy (kendricklamar2014):

Those are the steps to solve this Equation :)

OpenStudy (anonymous):

awesome thanks guys!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (kendricklamar2014):

You're Welcome :)